Please help me solve this problem. A student has 2 loans totalling $12,000. One of the loans is at 6% and the other is at 9%. If the total interest for both loans for a year is $855. What is the amount of each loan?
Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

.09x + .06($12,000-x) = 855
x = (855 - 720)/.03, amount at 9% = $4500
amount at 6% is ($12000-$4500)

Answer by richwmiller(17219)

Total amount of money invested: $12000
x+y=12000,
Total yearly interest for the two accounts is: $855
0.06*x+0.09*y=855
x=12000-y
Substitute for x
0.06*(12000-y)+0.09*y=855
Multiply out
720-0.06*y+0.09*y=855
Combine like terms.
0.03*y=135
Isolate y
y=$4500.00 at 9%
x=12000-y
Calculate x
x=$7500.00 at 6%
Check
0.06*7500+0.09*4500=855
450+405=855
855=855
If this statement is TRUE and neither x nor y is negative then all is well
code int1